I am stumped on this one.
Solve the system by addition.
-4x + 5y = 5
x - 2y = -2
Could you help guide me on the right track? Thank you.

Answer by stanbon(75887)

Solve the system by addition.
-4x + 5y = 5
x - 2y = -2
-------------
If you solve it by substitution you proceed as follows:
Solve the 2nd equation for x to get x = 2y-2
-----------
Substitute that value into the 1st equation and solve for y, as follows:
-4(2y-2)+5y = 5
-8y+8 + 5y = 5
-3y = -3
y = 1
-------------
Substitute that into x = 2y-2 to solve for "x", as follows:
x = 2*1 -2
x = 0
===============
Solution for the system (0,1)
==================
Cheers,
Stan H.